Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O3)
Montanari - 15:00 Friday 09 August 2019 (46629) Print this report
Operator Report - Morning shift

06:19UTC Adjusting to tune the working point (#46622), finished at 06:30UTC

06:28UTC unlocked (1st plot)
06:50UTC locked and in Science, but the ITF unlocked after few minutes (06:54UTC, 2nd plot)
07:14UTC locked and in Science

From 07:22UTC, ISC and SQZ had worked to improve sensibility and stability. Activity still in progress.

Air Conditioning
NE Chiller Failure. No effect in building's temperature.
A dedicated entry will be explain the problem.

Images attached to this report
Search Help
×

Warning

×